the probability that a flight arrives on time is 0.7 and the probability that a flight departs on time is 0.8.The probability that a flight neither arrives on time nor departs on time is 0.15. Then the probability that flight arrives on time and departs on time is?
1
Expert's answer
2020-09-07T12:45:55-0400

Here given,


 probability that a flight arrives on time = P(A) = 0.7

probability that a flight departs on time = P(B) = 0.8.

probability that a flight neither arrives on time nor departs on time = P(A' "\bigcap" B') = 0.15

probability that flight arrives on time and departs on time =P(A "\bigcap" B)


Using De'morgan law P(A "\bigcup" B)' = P(A' "\bigcap" B')

so, P(A "\bigcup" B)' = P(A' "\bigcap" B') = 0.15


P(A "\bigcup" B)' = 1- P(A "\bigcup" B)

P(A "\bigcup" B) = 1-P(A "\bigcup" B)'

=1-0.15

=0.85

Now,

P(A "\bigcup" B) = P(A) + P(B) -P(A "\bigcap" B)

P(A "\bigcap" B) = P(A) + P(B) - P(A "\bigcup" B)

= 0.7 + 0.8 - 0.85

P(A "\bigcap" B) = 0.65


ANSWER : probability that flight arrives on time and departs on time =P(A "\bigcap" B)=0.65